OpenStudy (anonymous):

What is the standard deviation of this data to the nearest one hundredth? 10 12, 8, 2

OpenStudy (amistre64):

whats the mean?

OpenStudy (anonymous):

8

OpenStudy (amistre64):

subtract 8 from every data point .. whats the set we get?

OpenStudy (anonymous):

2, 4, 0, -6

OpenStudy (amistre64):

good, now square those values

OpenStudy (anonymous):

How?

OpenStudy (amistre64):

n^2 = n*n

OpenStudy (anonymous):

Multiply each number with itself

OpenStudy (amistre64):

the longest part is simply finding the sum of the squares of the differences from the mean ....

OpenStudy (amistre64):

step 1; find the mean of the set step 2; subtract the mean from the set step 3; square the results step 4; find the mean of the squares (if its a population), or use n-1 for a sample the result of this is called variance. the sqrt(variance) equals standard deviation

OpenStudy (anonymous):

Ok i'm confused on the last step

OpenStudy (amistre64):

2, 4, 0, -6 ; square these values 2^2, 4^2, 0^2, (-6)^2 ; what do we get to work with?

OpenStudy (anonymous):

4, 16, 0, -36

OpenStudy (amistre64):

-6 * -6 = 36 the idea is to have all positive values now considering that the data set is a population, and not a sample of a population .... what is the mean of the squares?

OpenStudy (anonymous):

14

OpenStudy (amistre64):

then the variance is 14 standard deviation is the sqrt of variance soooo sqrt(14) = ??? rnd it as they desire

OpenStudy (anonymous):

Ok thanks

OpenStudy (amistre64):

suppose this was a sample from a given population .... such that the population had 32 data points. that mean of the squares would be refering to a sample (of size 4) of the population and we would need to divide by: 4-1 but thats just insight into a different question

OpenStudy (anonymous):

How do you round 14 to the nearest hundreth then?

OpenStudy (anonymous):

*one hundreth

OpenStudy (amistre64):

what do you get as the sqrt(14) ?

OpenStudy (anonymous):

14?

OpenStudy (anonymous):

or 196?

OpenStudy (amistre64):

\[\sqrt{14}=??\]

OpenStudy (amistre64):

im assuming you know how to operate a calculator for this

OpenStudy (amistre64):

if not, type: sqrt(14) into a google search

OpenStudy (anonymous):

3.741?

OpenStudy (amistre64):

yep, the "hundreths" spot is 2 decimals

OpenStudy (amistre64):

so, 3.74 is fine

OpenStudy (anonymous):

ok thanks

OpenStudy (amistre64):

youre welcome